OpenStudy (anonymous):

What is an equation of the line that passes through the point (-4,-1) and is parallel to and perpendicular to the line 2x+7y=14 ?

OpenStudy (anonymous):

i need help asap !

OpenStudy (anonymous):

First, change 2x + 7y into y=mx+b form.

OpenStudy (anonymous):

= 14*

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

First, move the "2x" to the other side by subtracting it to both sides, then divide by 7.

OpenStudy (anonymous):

What do you get? y = 2/7x + 7 Now, you need to find the perpendicular, so the new line must intersect at a 90 degree angle all-around (meaning all angles to the point equal 90 to make 360).

OpenStudy (anonymous):

Are you following this?

OpenStudy (anonymous):

sort of...

OpenStudy (anonymous):

Need a breakdown?

OpenStudy (anonymous):

yea. i just really need to get the answer asap

OpenStudy (anonymous):

If that's the case, just draw a line through the line i just gave you from the point (-4, -1)

OpenStudy (anonymous):

but i need equations

OpenStudy (anonymous):

Hold on, then.

OpenStudy (anonymous):

I got y = -1 2/7x - 1/7, but i'm not sure.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

Tell me if its right, kay?

OpenStudy (anonymous):

i dont think so cause i need 2 seperate equations. one for parallel and one for perpendicular...

OpenStudy (anonymous):

Then that would be the perpendicular. For the m, it should be 2/7x.